OpenStudy (anonymous):

The speed of a pendulum is 0.91 C and its time period is 1.8 sec. Then what will be its period in initial frame. A. 2 s B. 1.62 s C. 1.60 s D. 1.58 s

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

my goodness are you dealing with special with relativity?

OpenStudy (anonymous):

special relativity

OpenStudy (anonymous):

No idea :(

OpenStudy (anonymous):

@AllTehMaffs

OpenStudy (anonymous):

what do u mean no idea? u just typed 0.91 C.. i hate 0.91Cs C = speed of light right?

OpenStudy (anonymous):

Ops Sorry . I though u r asking Anyways yeah its C

OpenStudy (anonymous):

Speed of light

OpenStudy (anonymous):

so you are taking course on special relativity?!

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

hooray relativistic maths!

OpenStudy (anonymous):

not my forte.. !! see you :P

OpenStudy (anonymous):

its ok @Mashy

OpenStudy (anonymous):

\[\Delta T' = DeltaT / \sqrt{1-v ^{2}/c ^{2}}\]

OpenStudy (anonymous):

Am I right @AllTehMaffs ?

OpenStudy (anonymous):

In the inertial frame its time will be dilated, so its time should be longer methinks - 2's the only answer that's bigger. We can check though! \[ \Delta t_{inertial} = \gamma \Delta t_{pendulum}\] which is your t' and t?

OpenStudy (anonymous):

\[ \gamma=\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}\]

OpenStudy (anonymous):

t' is time dilated and t is time at rest

OpenStudy (anonymous):

And time "at rest" is the time in the frame of the pendulum, yah?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then yooz right! although I'm getting the wrong answer with it :/

OpenStudy (anonymous):

which one is correct ?

OpenStudy (anonymous):

oo.. its that simple? :D.. i thought that equation only worked for relativistic masses :D

OpenStudy (anonymous):

time dilation and length contraction, too ^_^ All derived from the same 4-space vector

OpenStudy (anonymous):

err, opposite for length contraction

OpenStudy (anonymous):

oh i ll be darned!

OpenStudy (anonymous):

I still dunno why the answer it's giving isn't right though :/ Are you sure it was .91c?

OpenStudy (anonymous):

we have to find period in initial frame . Is it Delta T ?

OpenStudy (anonymous):

.74 sec is mine answer ..

OpenStudy (anonymous):

I got 4.3 sec....

OpenStudy (anonymous):

t = time observed in the other reference frame t0 = time in observers own frame of reference (rest time) v = the speed of the moving object c = the speed of light in a vacuum

OpenStudy (anonymous):

yahyah, and the speeds are all given in the frame of the pendulum, right? I mean, even if they weren't, there's no combination of those numbers that gives any of those answers...

OpenStudy (anonymous):